In triangle ABC if B is 30 and c is root b then A can be equal to

Grade:9

1 Answers

Arun
6 years ago
Dear Rohan
 
\text{HINT: Use }\cos30^\circ=\dfrac{c^2+a^2-b^2}{2ca}=\dfrac{(\sqrt3b)^2+a^2-b^2)}{2\cdot\sqrt3b\cdot a}
 
 
 
Hope Now you can solve this. If any problem occurs please feel free to ask.
Regards
Arun (askIITians forum expert)